OpenStudy (anonymous):

Solved the system by substitution.if the system is inconsistent and has no solution,state this .if the system is dependent,write the form of the solution for any real number x. (1) x=4y+3 2x+5y=-7

OpenStudy (shamil98):

Substitute the value of x into the other equation. if the solution is all real numbers then it is dependent. A system of equations that has at least one solution is Consistent a system of equations that has no solution is inconsistent If it has 1 or no solution, it is independent

OpenStudy (shamil98):

2(4y+3) + 5y = -7 8y + 6 + 5y = -7 13y = -13 y= -1 x = 4(-1) + 3 x = -1 the solution is (-1,-1), and only one solution and is independent and consistent.

OpenStudy (anonymous):

And what about this problem : 2x-y=-15 5x+6y=5

OpenStudy (anonymous):

I solved it and this is what I got but not sure of my answer :(2.4,31)